OpenStudy (anonymous):

A solid is a hemisphere with 90 degree slice removed. r=8 in. volume=

OpenStudy (anonymous):

Well I'm assuming by a 90 degree slice that a quarter of it is taken out. So: V=(4/3)pi r^3. V=(3/4)(4/3)pi r^3 V=pi r^3 V=pi 8^3 V=512 pi in^3